The length of a rectangle is 3more than its width. the perimeter of the rectangle is 58cm. what are the rectangles dimension?
Answer by mananth(16946)

width =x
length = x+3
..
2*(l+w)=perimeter
2*(x+x+3)=58
2*(2x+3)=58
/2
2x+3=29
2x=29-3
2x=26
/2
x=13 the width
length = x+3====>13+3=16